# Towards Browser Controls to Protect Cookies from Malicious Extensions

### v2

@@ Abstract
Cookies maintain state across related web traffic. As such, cookies are commonly used for authentication by storing a user’s session ID and replacing the need to re-enter credentials in subsequent traffic. These so-called “session cookies” are prime targets for attacks that aim to steal them to gain unauthorized access to user accounts. To mitigate these attacks, the Secure and HttpOnly cookie attributes limit a cookie’s accessibility from malicious networks and websites. However, these controls overlook browser extensions: third-party HTML/JavaScript add-ons with access to privileged browser APIs and the ability to operate across multiple websites. Thus malicious or compromised extensions can provide unrestricted access to a user’s session cookies.
@@
In this work, we first analyze the prevalence of extensions with access to “risky” APIs (those that enable cookie modification and theft) and find that they have hundreds of millions of users. Motivated by this, we propose a mechanism to protect cookies from malicious extensions by introducing two new cookie attributes: BrowserOnly and Monitored. The BrowserOnly attribute prevents extension access to cookies altogether. While effective, not all cookies can be made inaccessible. Thus cookies with the Monitored attribute remain accessible but are tied to a single browser and any changes made to the cookie are logged. As a result, stolen Monitored cookies are unusable outside their original browser and servers can validate the modifications performed. To demonstrate the proposed functionalities, we design and implement CREAM (Cookie Restrictions for Extension Abuse Mitigation) a modified version of the open-source Chromium browser realizing these controls. Our evaluation indicates that CREAM effectively protects cookies from malicious extensions while incurring little run-time overheads.
1 Introduction
@@
Cookies remain a universal state management mechanism on the web. They allow servers to store data on the user’s browser and for that data to be automatically returned in all subsequent requests [1]. This allows servers to link related traffic into a single session and maintain data across the session.
@@
A common and security-critical use of cookies is user authentication, where cookies store an authenticated user session ID and act as a replacement for credentials [2, 3]. Naturally, this makes cookies a valuable target for attacks such as session hijacking, session fixation, and Cross-Site Request Forgery (CSRF). Session hijacking [4, 5, 2] attempts to gain access to the user’s account by stealing the value of their session cookie. These attacks are not limited to session cookies alone as non-session cookies can still leak users’ private information [6]. Similarly, session fixation [3, 5, 7] injects a malicious session cookie with a pre-known value into the user’s browser. Once the user authenticates with the website, an adversary can use the pre-known ID to access the user’s account. Finally, CSRF tricks the browser into issuing a cross-site request to perform some action as the user without their knowledge [3, 8, 5].
@@
Several cookie features have been introduced to combat these attacks [9, 1]. The Secure attribute restricts cookies to HTTPS traffic. Thus Secure cookies are always encrypted on the network. Similarly, the HttpOnly attribute prevents JavaScript (JS) from accessing cookies. Together these attributes help mitigate session hijacking by preventing cookie theft from malicious JS and over the network. The SameSite attribute restricts how cookies are attached to cross-site requests, reducing the threat of CSRF. Cookies also define two cookie prefixes: __Secure- and __Host-. These prefixes enforce that certain criteria be true when creating the cookie. As such, cookie prefixes help reduce the threat of session fixation by asserting how a cookie was created.
@@
These controls reduce the threat posed by malicious websites and networks, however, browsers also support third-party HTM/JS applications called extensions [10, 11]. Extensions add functionality to browsers, run across many websites, and can bypass several browser controls by default, including cookie controls [12, 13]. Despite being JS themselves, extensions are not restricted by HttpOnly. Similarly, extensions can access network traffic as it is processed in the browser. Due to this, they can access cookies within network headers before TLS encryption, bypassing the Secure attribute. Further, extensions can add cookies to traffic, ignoring the SameSite attribute. Extensions can also create cookies with known prefixes and edit cookies/remove their security attributes. As such, extensions still pose a large threat to cookies.
@@
These permissive extension controls stem from the current browser trust model. Existing browser controls often assume two possible attackers, a web attacker [14, 15, 16, 17, 18, 19, 20, 21] and a network attacker [3, 15, 16, 18, 19]. In both models, the attacker controls a malicious website, whereas a network attacker can also inspect and tamper with HTTP traffic. In both models, the attacker is an external party. Thus, everything inside the browser is considered trusted, including extensions. This is also reflected by the extension architecture itself assuming “benign but buggy" extensions [19].
@@
Instead of architectural controls, browser vendors try to prevent malicious extensions from becoming available in the first place. Extensions are reviewed to adhere to a browser’s extension policies [22, 23] before being listed publicly. However, malicious extensions can pass this review [24, 25, 26], requiring them to be pulled from browsers after the fact [27, 28]. Extension permissions must also be granted at installation [29]. Thus users can prevent malicious extensions by denying them permission, however, users are often unaware of the risks associated with these permissions. [30]. Regardless of either check, extensions can also "become" malicious. Extensions update automatically in the background [31]. As such, extensions can silently receive a malicious update while maintaining their permissions [26]. Browsers also allow the manual installation of extensions for development [32]. This can be leveraged to bypass all checks by tricking the user into manually installing a malicious extension.
@@
Prior work also demonstrates that vulnerable extensions can leak their permissions to a malicious page [33, 18, 34]. Thus, compromised extensions can also be used to bypass existing cookie controls. To demonstrate this risk, we use the extension APIs to hijack test accounts on several websites (see Section 3). We also scrape 156,783 extensions from both Chrome’s and Firefox’s extension repositories to assess the prevalence of these APIs in current extensions.
@@
In this work, we aim to reevaluate the trust model employed by current browsers to mitigate risks posed by extensions. Prior work has sought to protect sessions from extensions by preventing their access to sensitive information on web pages [35, 36, 37]. While these controls protect sessions from direct access by extensions, malicious extensions maintain access to session cookies. Thus adversaries can still gain unauthorized access to the session through these cookies.
@@
We propose two new cookie attributes to protect cookies from malicious extensions: BrowserOnly and Monitored. Unlike previous controls, we assume extensions are untrusted and potentially malicious. The BrowserOnly attribute prevents extensions from accessing cookies outright. However, some cookies must remain accessible for websites and extensions to function properly. Thus the Monitored attribute allows extensions to access cookies, ties cookies to a single browser instance, and maintains a changelog for each cookie. This changelog and other server-defined data are authenticated and sent to the server alongside each cookie. While Monitored does not prevent cookie theft, it enables server-side detection of stolen and illegally modified cookies. Note that these features do not imply privacy loss (see Section 8).
@@
The Monitored attribute uses symmetric keys to link cookies to a single browser and authenticate their changelog-s. Specifically, the browser generates a unique key for each site it visits and attaches the key to all outbound HTTPS requests to its respective site. The server then uses the received key to identify the browser by saving the key in the cookie as part of the Monitored attribute. When the server receives the authenticated attribute in later requests, it can verify it using the saved key. This verification ensures that the cookie came from the correct browser and its changelog has not been tampered with. While these keys allow individual sites to identify the user’s browser, they cannot be used to track the user across different sites as the browser uses a unique key for each site.
@@
Support for the BrowserOnly and Monitored attributes is obtained by modifying the browser architecture to revoke an extension’s ability to access/modify these attributes before their usage by the underlying browser core. As a proof of concept, we implement these features as an open-source modified version of the Chromium browser [38] and evaluate our prototype concerning storage and run-time impacts. In summary, our intended contributions are as follows:
@@
    We study Chrome’s and Firefox’s extension repositories to assess the prevalence of extensions that, if exploited, could threaten session cookies.
@@
    We propose a new browser trust model regarding extensions, in which we introduce two new cookie attributes to protect cookies from potentially malicious browser extensions. The BrowserOnly attribute prevents cookies from being accessed by extensions outright. The Monitored attribute leaves cookies accessible while allowing servers to detect cookie theft and tampering (allowing servers to refuse service to stolen cookies). To support this functionality, we design a mechanism to tie browser instances to cookies without linking users across different domains.
@@
    To realize this concept we implement and open-source [39] CREAM (Cookie Restrictions for Extension Abuse Mitigation) as a modified version of Chromium and evaluate its functionality and associated overheads.
2 Background
2.1 Browser Architecture
@@
Renderer processes parse untrusted, potentially malicious, web content (i.e. HTML, CSS, JS) and build the final page displayed to the user [43, 17, 45]. As such, they are unprivileged. Renderers also execute in a sandbox to further restrict their access to the underlying host. This forces all host-related requests to go through the browser process.
@@
Browsers are also service-oriented [46, 47]. Many renderer and browser process features have been moved to independent services that are often executed within separate processes. For example, browsers have a Network Service that implements the browser’s network stack (all low-level networking code) and handles all network transactions [48, 42, 49]. As these feature-specific processes have no standard name, we use Chrome’s nomenclature of “utility processes” [43, 44] to refer to them. Utility processes can be privileged, unprivileged, sandboxed, or unsandboxed depending on their functionality.
In the remainder of this paper, we often reference “the browser”. When referring to “the browser” we are discussing the collection of features commonly referred to as the browser. This includes both the privileged browser process as well as utility processes, such as the Network Service. We refer to the browser process alone as the “browser core” instead.
@@
2.2 Extension Architecture
All extensions have a manifest.json file that describes them [11, 19]. Currently, there are two manifest versions (V2 and V3), with slight differences in how attributes are defined and which extension APIs are available. This file lists the APIs and scripts an extension uses, websites it runs on, how it is loaded, resources it can access, and more. The browser requires the manifest to be present when an extension is installed and restricts the extension’s behavior to what is stated in its manifest. This prevents extensions from gaining more privileges after installation, except for optional permissions which can be requested at any time, but must still be declared in the manifest file at installation.
@@ 2.2 Extension Architecture
This isolation prevents the extension core from interacting directly with any website. Instead, content scripts are injected into the page and have direct access to its content [19, 18]. These scripts execute within the renderer process of the website and are thus at an increased risk of attack. Due to this, content scripts are isolated from other scripts on the page [19, 33]. Similarly, content scripts are unprivileged and cannot access the extension APIs. Instead, content scripts can only send messages to the extension core using the message-passing API. This aims to prevent a compromised renderer from gaining the powers of any extensions embedded within them.
2.3 Cookies & Cookie Management
At their simplest, cookies are key-value pairs associated with the server that creates them. Cookies can also set multiple optional attributes that further define how they are handled. Which server a cookie belongs to and what requests it should be attached to are determined by the Domain and Path attributes [1]. These indicate what websites and which portion of those sites the cookie can be sent to respectively. If not specified by the server, the browser uses the hostname and path of the website that created the cookie by default. Cookies can also be partitioned. Partitioning limits a cookie’s accessibility to its embedding context such that cookies set by the same site but embedded on different pages are not accessible to each other [51]. This is often used to prevent third-party cookies from tracking users across websites. As discussed in Section 1, cookies also have three security attributes: HttpOnly, Secure, and SameSite. They also support two prefixes that enforce certain criteria when creating a cookie [9].
2.3.1 Cookie Manager
Cookies can be set in many ways. Most commonly, they are defined in the Set-Cookie HTTP header [52] and created by the Network Service [53]. Since cookies provide state for network requests, they are stored within and are directly interactable from the Network Service. Cookies can also be created from a webpage through ‘document.cookie’ [54] or the Cookie Store API [55], by the browser core, and by extensions with the Cookies extension API permission [56]. However, none of these can interact with cookies directly. Instead, the Network Service exposes APIs for accessing its internal cookie store [53]. The arrows in Figure 1 depict how different browser components access cookies.
@@ 2.3.1 Cookie Manager
2.3.2 Cookie Monster
All cookies are maintained by the browser’s cookie store. Chromium’s concrete implementation of the cookie store is called the Cookie Monster [57]. There are two ways to fetch cookies from the Cookie Monster [58]. Either the Cookie Monster can return all unexpired cookies or a subset of cookies that match some provided criteria. We refer to these methods as GetAll and GetList in the rest of this paper. When using GetAll, the Cookie Monster returns all cookies within its store, regardless of their attributes. However, most fetches use GetList which takes a domain and a set of options, fetches all the provided domain’s cookies, and filters them based on the provided options [53]. These options include checks such as whether to allow HttpOnly or Secure cookies in the results.
To set a cookie (Set operation), the caller (browser core, renderer, extension, etc.) creates a new cookie and sends it to the Cookie Monster along with its domain and other cookie options [58]. Upon receipt of the cookie, the Cookie Monster first checks that the cookie is valid. Similar to GetList, this validation uses the supplied options to ensure the cookie is properly configured (e.g. verifying that an HttpOnly cookie is not being set by JS). If valid, the Cookie Monster then checks for any existing matching cookie. If a match is found, it is deleted and replaced with the new cookie. Otherwise, the cookie is added to the internal map of cookies as a new entry.
3 On the Extent of Extension Threats
3.1 Manipulating Cookies via Extension APIs
To illustrate the practical threats extensions pose to cookies, we implemented three test extensions targeting different extension APIs and tested them in Chrome. Each test extension was built in both manifest versions (V2 and V3), when applicable. We specifically target the Cookies, WebRequest, and DeclarativeNetRequest APIs, as they allow interaction with cookies. The Cookies API allows extensions to get, set, remove, and be notified of changes to cookies within the browser [56]. The cookies accessible to an extension are limited to those of websites it has permission to run on. Nonetheless, an extension can have permission to run on all websites (which is not uncommon). The WebRequest API allows extensions to view network traffic mid-processing [59]. As such, it allows extensions to access HTTP headers, including cookie headers. The WebRequest API can also be extended with the WebRequestBlocking permission to allow extensions to pause the execution of network events and edit request headers. Finally, the DeclarativeNetRequest API also allows extensions to modify network requests [60, 61]. However, unlike WebRequest, DeclarativeNetRequest cannot view the contents of the request. Each test extension targeted a single API as follows:

### v3

@@ Abstract
Cookies are vital to the web’s current operation, taking on several important and security-critical tasks such as user authentication. Due to their importance, browsers implement several cookie controls, such as the Secure and HttpOnly attributes to help protect cookies from malicious networks and websites. While effective, these controls overlook browser extensions: third-party HTML/JavaScript add-ons with access to privileged browser APIs and the ability to operate across multiple websites. As a result, malicious/compromised extensions can gain unrestricted access to a user’s cookies.
@@
In this work, we first analyze the prevalence of extensions that can access and modify cookies. Our analysis shows that these extensions reach hundreds of millions of users. We then propose a mechanism to protect cookies from malicious (or compromised) extensions, introducing two complementary cookie attributes: BrowserOnly and Monitored. The BrowserOnly attribute prevents extensions from accessing a cookie outright. However, not all cookies can be made inaccessible while preserving intended functionality. As such, the Monitored attribute serves as an alternative, maintaining accessibility while tying cookies to a single browser instance and logging any modifications made to the cookie by extensions. As a result, stolen Monitored cookies are unusable outside their original browser and servers can verify the logged modifications. To demonstrate the proposed concepts, we design and implement CREAM (Cookie Restrictions for Extension Abuse Mitigation) as a modified version of Chromium realizing these controls. Our evaluation shows effective cookie protection against malicious extensions and low run-time overheads.
I Introduction
@@
Cookies remain a universal state management mechanism on the web. They allow servers to store data on the user’s browser and for that data to be automatically returned in all subsequent requests [1]. This enables the linking of related traffic into a session and the maintenance of data across it.
@@
A common and security-critical use of cookies is user authentication, where cookies store an authenticated user session ID and replace credentials in future traffic [2, 3]. This makes cookies a valuable target for attacks such as session hijacking, session fixation, and Cross-Site Request Forgery (CSRF). Session hijacking [4, 5, 2] attempts to gain access to the user’s account by stealing the value of their “session cookie”. These attacks are not limited to session cookies alone as non-session cookies can still leak users’ private information [6]. Similarly, session fixation [3, 5, 7] injects a malicious session cookie with a known value into the user’s browser. Once the user authenticates with the website, an adversary can use the known value to access the user’s account. CSRF tricks the browser into issuing a cross-site request to perform some action as the user without their knowledge [3, 8, 5].
@@
Several cookie features have been introduced to combat these attacks [9, 1]. The Secure attribute restricts cookies to HTTPS traffic, ensuring they are always encrypted on the network. Similarly, the HttpOnly attribute prevents JavaScript (JS) from accessing cookies. Together, these attributes help mitigate session hijacking by preventing cookie theft from malicious JS and networks. The SameSite attribute restricts how cookies are attached to cross-site requests, reducing the threat of CSRF. Cookies also define two cookie prefixes: __Secure- and __Host-. These prefixes enforce that certain criteria are true when creating the cookie. As such, cookie prefixes help reduce the threat of session fixation by asserting how a cookie was created.
@@
While effective against malicious websites and networks, the aforementioned controls do not protect against browser extensions. Extensions are third-party HTML/JS applications that add functionality to browsers and execute across multiple websites [10, 11]. As such, extensions can bypass several browser controls by default, including cookie controls [12, 13]. For example, extensions can access HTTPOnly cookies despite being JS themselves. Similarly, extensions can access network traffic as it is processed in the browser. Thus, they can view and modify cookies within network headers despite TLS encryption, bypassing both the Secure and SameSite attributes. Further, in the current model, extensions can simply modify cookies to remove their security attributes altogether. Unsurprisingly, malicious/compromised extensions have been widely used in recent attacks [14, 15, 16] to steal cookies and gain access to user-authenticated sessions.
@@
These permissive extension controls stem from the current browser trust model. Most existing browser controls assume two possible attackers, a web attacker [17, 18, 19, 20, 21, 22, 23, 24] and a network attacker [3, 18, 19, 21, 22]. In both models, the attacker controls a malicious website, whereas a network attacker can also inspect and tamper with HTTP traffic. In both models, the attacker is an external party. Thus, everything inside the browser is considered trusted, including extensions. This is also reflected by the extension architecture itself assuming “benign but buggy” extensions [22].
@@
Instead of architectural controls, browser vendors try to prevent malicious extensions from becoming available in the first place. Before being publicly listed, extensions are reviewed to adhere to the browser’s extension policies [25, 26]. However, malicious extensions can still pass this review [27, 28, 29, 30], requiring them to be pulled from users’ browsers after the fact [31, 32]. Along with this, most extension permissions must be accepted at installation [33]. Thus users can prevent malicious extension behavior by denying them permissions. However, users are often unaware of the risks associated with these permissions [34]. Regardless of either check, extensions can also “become” malicious. Browsers automatically update extensions in the background [35]. As such, extensions can silently receive a malicious update while maintaining their current permissions [29]. Further, browsers allow installing extensions manually (typically for extension development) [36]. This can be leveraged to bypass all checks by tricking the user into manually installing a malicious extension.
@@
Prior work also demonstrates that vulnerable extensions can leak their permissions to a malicious website [37, 21, 38]. Thus, compromised extensions can also be used to bypass existing cookie controls. To demonstrate this risk, we use the extension APIs to hijack test accounts on several websites (see Section III). We also scrape 156,783 extensions from both Chrome’s and Firefox’s extension repositories to assess the prevalence of these APIs in current extensions.
@@
Motivated by these issues, this work aims to reevaluate the trust model employed by current browsers to mitigate risks posed by extensions. Prior work has sought to protect sessions from extensions by preventing their access to sensitive information on web pages [39, 40, 41]. While these controls protect sessions from direct access by extensions, malicious extensions maintain access to session cookies. Thus adversaries can still gain unauthorized access to the session.
Contributions.
@@
We propose two new cookie attributes to protect cookies from malicious extensions: BrowserOnly and Monitored. Unlike previous controls, we consider untrusted and potentially malicious extensions (e.g., as in recent attacks of [14, 15, 16]). The BrowserOnly attribute prevents extensions from accessing cookies. While effective, blocking access to certain cookies would break some extension and website functionality. For these cases, the Monitored attribute allows extensions access to cookies but ties the cookies to a single browser instance and maintains a changelog for each cookie. When sending a Monitored cookie to a server, the changelog and other server-defined data is authenticated and sent alongside the cookie. As such, Monitored does not prevent direct cookie theft. Instead, it enables server-side detection of stolen and illegally modified cookies, preventing their subsequent use. Importantly, as discussed in Section VIII, the proposed features do not imply privacy loss.
@@
The Monitored attribute uses symmetric keys to link cookies to a single browser and authenticate their chagelog-s. Specifically, the browser generates a unique key for each site it visits and attaches the key to all outbound HTTPS requests to its respective site. The server then embeds this key as part of a cookie’s Monitored attribute to identify the browser and verify the authenticated attribute in future requests. This verification ensures the cookie came from the correct browser and its changelog has not been tampered with. While these keys allow individual sites to identify the user’s browser, they cannot be used for cross-site tracking as the browser uses a unique key for each site visited.
@@
Enforcement of the BrowserOnly and Monitored attributes is obtained by modifying the browser architecture to revoke extensions’ ability to access/modify these attributes before the underlying browser core. As a proof of concept, we implement these features as an open-source modified version of the Chromium browser [42] and evaluate our prototype concerning storage and run-time impacts. In summary, our intended contributions are:
@@
    We study Chrome’s and Firefox’s extension repositories to assess the prevalence of extensions that, if malicious or compromised, could threaten cookies.
@@
    We propose a new browser trust model regarding extensions, in which we introduce two new cookie attributes to protect cookies from potentially malicious browser extensions. The BrowserOnly attribute prevents cookies from being accessed by extensions outright. The Monitored attribute leaves cookies accessible while allowing servers to detect cookie theft and tampering (allowing servers to refuse service to stolen cookies). To support this functionality, we design a mechanism to tie browser instances to cookies without linking users across domains.
@@
    To realize this concept we implement and open-source CREAM [43] (Cookie Restrictions for Extension Abuse Mitigation) as a modified version of Chromium and evaluate its functionality and (modest) overheads.
II Background
II-A Browser Architecture
@@
Renderer processes parse untrusted, potentially malicious, web content (i.e., HTML, CSS, JS) and build the final page displayed to the user [47, 20, 49]. As such, they are unprivileged. Renderers also execute in a sandbox to further restrict their access to the underlying host. This forces all host-related requests to go through the browser process.
@@
Browsers are also service-oriented [50, 51]. Many renderer and browser features have been moved to independent services that often execute within separate processes. For example, the browser’s Network Service implements its network stack (all low-level networking code) and handles all network transactions [52, 46, 53]. As these feature-specific processes have no standard name, we use Chrome’s nomenclature of “utility processes” [47, 48] to refer to them. Utility processes can be privileged, unprivileged, sandboxed, or unsandboxed depending on their functionality.
In the remainder of this paper, when referencing “the browser” we are discussing the collection of features commonly referred to as the browser. This includes both the privileged browser process and utility processes, such as the Network Service. We refer to the browser process alone as the “browser core” instead.
@@
II-B Extension Architecture
All extensions have a manifest.json file [11, 22]. Currently, there are two manifest versions (V2 and V3), with slight differences in how attributes are defined, which extension APIs are available, and which browsers support them. This file lists the APIs and scripts an extension uses, websites it runs on, how it is loaded, resources it can access, and more. The browser requires the manifest to be present when an extension is installed and restricts the extension’s behavior to what is stated in its manifest. This prevents extensions from gaining more privileges after installation, except for optional permissions which can be requested at any time, but must still be declared in the manifest file at installation.
@@ II-B Extension Architecture
This isolation prevents the extension core from interacting directly with any website. Instead, content scripts are injected into the page and have direct access to its content [22, 21]. These scripts execute within the renderer process of the website and are thus at an increased risk of attack. Due to this, content scripts are isolated from other scripts on the page [22, 37]. Similarly, content scripts are unprivileged and cannot access most extension APIs. Instead, content scripts can only send messages to the extension core using the message-passing API. This aims to prevent a compromised renderer from gaining the powers of any extensions embedded within it.
II-C Cookies & Cookie Management
At their simplest, cookies are key-value pairs associated with the server that creates them. Cookies can also set multiple optional attributes that further define how they are handled. Which server a cookie belongs to and what requests it should be attached to are determined by the Domain and Path attributes [1]. These indicate which websites and which portion of those sites the cookie can be sent to respectively. If not specified, the browser uses the hostname and path of the website that created the cookie by default. Cookies can also be partitioned. Partitioning limits a cookie’s accessibility to its embedding context such that cookies set by the same site but embedded on different pages are not accessible to each other [55]. This is often used to prevent third-party cookies from tracking users across websites. As discussed in Section I, cookies also have three security attributes: HttpOnly, Secure, and SameSite. They also support two prefixes that enforce certain cookie creation criteria [9].
II-C1 Cookie Manager
Cookies can be set in many ways. Most commonly, they are defined in the Set-Cookie HTTP header [56] and created by the Network Service [57]. As cookies provide state for network requests, they are stored within and are directly interactable from the Network Service. Cookies can also be created from a webpage through ‘document.cookie’ [58] or the Cookie Store API [59], by the browser core, and by extensions through the Cookies extension API [60]. However, none of these can interact with cookies directly. Instead, the Network Service exposes APIs for accessing its internal cookie store [57]. The arrows in Figure 1 depict how different browser components access cookies.
@@ II-C1 Cookie Manager
II-C2 Cookie Monster
All cookies are maintained by the browser’s cookie store. Chromium’s concrete implementation of the cookie store is called the Cookie Monster [61]. There are two ways to fetch cookies from the Cookie Monster [62]. The Cookie Monster can either return all unexpired cookies or a subset of cookies that match some provided criteria. We refer to these methods as GetAll and GetList respectively. When using GetAll, the Cookie Monster returns all cookies within its store, regardless of their attributes. However, most fetches use GetList which takes a domain and a set of options, fetches all the provided domain’s cookies, and filters them based on the provided options [57]. These options include checks such as whether to allow HttpOnly or Secure cookies in the results.
To set a cookie (Set operation), the caller (browser core, renderer, extension, etc.) creates a new cookie and sends it to the Cookie Monster along with its domain and other cookie options [62]. Upon receipt of the cookie, the Cookie Monster first checks that the cookie is valid. Similar to GetList, this validation uses the supplied options to ensure the cookie is properly configured (e.g., verifying JS is not setting an HTTPOnly cookie). If valid, the Cookie Monster then checks for any existing matching cookie. If a match is found, it is replaced with the new cookie. Otherwise, the cookie is added to the internal map of cookies as a new entry.
III On the Extent of Extension Threats
III-A Manipulating Cookies via Extension APIs
To illustrate the practical threats extensions pose to cookies, we implemented three test extensions targeting different extension APIs. Each extension was tested in Chrome against test accounts on several websites. Each test extension was built with both V2 and V3 manifest versions, when applicable. V2 extensions were recently deprecated in Chrome [63], however, at the time of testing they were still supported and V2 maintains support in other browsers [64]. We specifically target the Cookies, WebRequest, and DeclarativeNetRequest APIs, as they allow interaction with cookies. The Cookies API allows extensions to get, set, remove, and be notified of changes to cookies within the browser [60]. The cookies accessible to an extension are limited to those of websites it has permission to run on. Nonetheless, an extension can have permission to run on all websites (which is not uncommon). The WebRequest API allows extensions to view network traffic mid-processing [65]. As such, it allows extensions to access HTTP headers, including cookie headers. The WebRequest API can also be extended with the WebRequestBlocking permission to allow extensions to pause the execution of network events and edit request headers as well. Finally, the DeclarativeNetRequest API also allows extensions to modify network requests [66, 67]. However, unlike WebRequest, DeclarativeNetRequest cannot view the contents of the request. Each test extension targeted a single API as follows: